GCC for the Atari ST

Using the GNU C-Compiler on the Atari ST by Frank Ridderbusch 1. Draft March 9, 1990 Copyright \begin{matrix}c \\ ◯ \end{matrix} 1988, 1989, 1990 Free Software Foundation, Inc.

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided also that the section entitled “GNU CC General Public License” is included exactly as in the original, and provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one.

Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that the section entitled “GNU CC General Public License” and this permission notice may be included in translations approved by the Free Software Foundation instead of in the original English.

GNU CC GENERAL PUBLIC LICENSE

(Clarified 11 Feb 1988)

The license agreements of most software companies keep you at the mercy of those companies. By contrast, our general public license is intended to give everyone the right to share GNU CC. To make sure that you get the rights we want you to have, we need to make restrictions that forbid anyone to deny you these rights or to ask you to surrender the rights. Hence this license agreement.

Specifically, we want to make sure that you have the right to give away copies of GNU CC, that you receive source code or else can get it if you want it, that you can change GNU CC or use pieces of it in new free programs, and that you know you can do these things.

To make sure that everyone has such rights, we have to forbid you to deprive anyone else of these rights. For example, if you distribute copies of GNU CC, you must give the recipients all the rights that you have. You must make sure that they, too, receive or can get the source code. And you must tell them their rights.

Also, for our own protection, we must make certain that everyone finds out that there is no warranty for GNU CC. If GNU CC is modified by someone else and passed on, we want its recipients to know that what they have is not what we distributed, so that any problems introduced by others will not reflect on our reputation.

Therefore we (Richard Stallman and the Free Software Foundation, Inc.) make the following terms which say what you must do to be allowed to distribute or change GNU CC.

COPYING POLICIES

1. You may copy and distribute verbatim copies of GNU CC source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy a valid copyright notice “Copyright \begin{matrix}c \\ ◯ \end{matrix} 1988 Free Software Foundation, Inc.” (or with whatever year is appropriate); keep intact the notices on all files that refer to this License Agreement and to the absence of any warranty; and give any other recipients of the GNU CC program a copy of this License Agreement along with the program. You may charge a distribution fee for the physical act of transferring a copy.

2. You may modify your copy or copies of GNU CC or any portion of it, and copy and distribute such modifications under the terms of Paragraph 1 above, provided that you also do the following:

• cause the modified files to carry prominent notices stating that you changed the files and the date of any change; and

• cause the whole of any work that you distribute or publish, that in whole or in part contains or is a derivative of GNU CC or any part thereof, to be licensed at no charge to all third parties on terms identical to those contained in this License Agreement (except that you may choose to grant more extensive warranty protection to some or all third parties, at your option).

• You may charge a distribution fee for the physical act of transferring a copy, and you may

at your option offer warranty protection in exchange for a fee. Mere aggregation of another unrelated program with this program (or its derivative) on a volume of a storage or distribution medium does not bring the other program under the scope of these terms.

3. You may copy and distribute GNU CC (or a portion or derivative of it, under Paragraph 2) in object code or executable form under the terms of Paragraphs 1 and 2 above provided that you also do one of the following:

• accompany it with the complete corresponding machine-readable source code, which must be distributed under the terms of Paragraphs 1 and 2 above; or,

• accompany it with a written offer, valid for at least three years, to give any third party free (except for a nominal shipping charge) a complete machine-readable copy of the corresponding source code, to be distributed under the terms of Paragraphs 1 and 2 above; or,

• accompany it with the information you received as to where the corresponding source code may be obtained. (This alternative is allowed only for noncommercial distribution and only if you received the program in object code or executable form alone.)

For an executable file, complete source code means all the source code for all modules it contains; but, as a special exception, it need not include source code for modules which are standard libraries that accompany the operating system on which the executable file runs.

4. You may not copy, sublicense, distribute or transfer GNU CC except as expressly provided under this License Agreement. Any attempt otherwise to copy, sublicense, distribute or transfer GNU CC is void and your rights to use the program under this License agreement shall be automatically terminated. However, parties who have received computer software programs from you with this License Agreement will not have their licenses terminated so long as such parties remain in full compliance.

5. If you wish to incorporate parts of GNU CC into other free programs whose distribution conditions are different, write to the Free Software Foundation at 675 Mass Ave, Cambridge, MA 02139. We have not yet worked out a simple rule that can be stated here, but we will often permit this. We will be guided by the two goals of preserving the free status of all derivatives of our free software and of promoting the sharing and reuse of software.

Your comments and suggestions about our licensing policies and our software are welcome! Please contact the Free Software Foundation, Inc., 675 Mass Ave, Cambridge, MA 02139, or call (617) 876-3296.

NO WARRANTY

BECAUSE GNU CC IS LICENSED FREE OF CHARGE, WE PROVIDE ABSOLUTELY NO WARRANTY, TO THE EXTENT PERMITTED BY APPLICABLE STATE LAW. EX-CEPT WHEN OTHERWISE STATED IN WRITING, FREE SOFTWARE FOUNDATION, INC, RICHARD M. STALLMAN AND/OR OTHER PARTIES PROVIDE GNU CC "AS IS" WITH-OUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PER-FORMANCE OF GNU CC IS WITH YOU. SHOULD GNU CC PROVE DEFECTIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW WILL RICHARD M. STALL-MAN, THE FREE SOFTWARE FOUNDATION, INC., AND/OR ANY OTHER PARTY WHO MAY MODIFY AND REDISTRIBUTE GNU CC AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY LOST PROFITS, LOST MONIES, OR OTHER SPE-CIAL, INCIDENTAL OR CONSEQUENTIAL DAMAGES ARISING OUT OF THE USE OR IN-ABILITY TO USE (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUSTAINED BY THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPERATE WITH ANY OTHER PROGRAMS) GNU CC, EVEN IF YOU HAVE BEEN ADVISED OF THE POSSIBILITY OF SUCH DAMAGES, OR FOR ANY CLAIM BY ANY OTHER PARTY.

Contributors to GNU CC

In addition to Richard Stallman, several people have written parts of GNU CC.

• The idea of using RTL and some of the optimization ideas came from the U. of Arizona Portable Optimizer, written by Jack Davidson and Christopher Fraser. See “Register Allocation and Exhaustive Peephole Optimization”, Software Practice and Experience 14 (9), Sept. 1984, 857-866.

• Paul Rubin wrote most of the preprocessor.

• Leonard Tower wrote parts of the parser, RTL generator, RTL definitions, and of the Vax machine description.

• Ted Lemon wrote parts of the RTL reader and printer.

• Jim Wilson implemented loop strength reduction and some other loop optimizations.

• Nobuyuki Hikichi of Software Research Associates, Tokyo, contributed the support for the SONY NEWS machine.

• Charles LaBrec contributed the support for the Integrated Solutions 68020 system.

• Michael Tiemann of MCC wrote most of the description of the National Semiconductor 32000 series cpu. He also wrote the code for inline function integration and for the SPARC cpu and Motorola 88000 cpu and part of the Sun FPA support.

• Jan Stein of the Chalmers Computer Society provided support for Genix, as well as part of the 32000 machine description.

• Randy Smith finished the Sun FPA support.

• Robert Brown implemented the support for Encore 32000 systems.

• David Kashtan of SRI adapted GNU CC to the Vomit-Making System.

• Alex Crain provided changes for the 3b1.

• Greg Satz and Chris Hanson assisted in making GNU CC work on HP-UX for the 9000 series 300.

• William Schelter did most of the work on the Intel 80386 support.

• Christopher Smith did the port for Convex machines.

• Paul Petersen wrote the machine description for the Alliant FX/8.

The following people contributed specially to the version for the Atari ST.

• John R. Dunning did the original port to the Atari ST.

• Jwahar R. Bammi improved the port. Jwahar Bammi and Eric Smith put together and maintain the current libraries. The Atari ST port has greatly benefited from contributions and Contributors to GNU CC 5

ideas of Edgar Roeder, Dale Schumacher, Kai–Uwe Bloem, Allan Pratt, John Dunning, Henry Spencer and many enthusiastic users in the atari community.

• Frank Ridderbusch compiled the manual for the Atari ST.

Introduction

This manual documents how to install and run the GNU C compiler on the Atari ST. It does not give an introduction in C or M68000 assembler. There is enough material on both subjects available. The user, who is familiar with a C compiler, that runs on a U**x system, should have no trouble at all to get GNU C running on the Atari ST. This manual was compiled from existing GNU manuals and various bits and pieces from John R. Dunning and Jwahar R. Bammi.

The sections, that describe the compiler driver and the preprocessor are nearly verbatim copies of sections in the respective manuals. The original manuals (Using and Porting GNU CC and The C Preprocessor), were written by Richard M. Stallmann. I modified these sections by removing material, which described features of GNU C for systems like Vaxen or Suns. To keep this manual resonably compact, I extracted only the sections, which describe the supported command options (and predefined macros in case of the preprocessor). If the user is interested in the extensions and details, which are implemented in GNU C, he has to refer to the original manuals. Whether all described options are usefull on the Atari has to be decided.

The facts, which are presented in the assembler and utility sections are mostly derived from the sources of the repective programs (from a cross compiler kit by J. R. Bammi based on GNU C 1.31), which were available to me. Other facts were gathered by try and error. So, these sections may be a bit shaky.

Please send any comments, corrections, etc concering this manual to

     Snail:
     Frank Ridderbusch
     Sander Str. 17
     4790 Paderborn, West Germany
     Email:
     BIX:  fridder
     UUCP: ridderbusch.kd@sni-usa.com

1 Installing GCC

The compressed archive of the GNU C compiler binary distribution contains the ’common’ executables of the GNU compiler. That means the compiler driver (‘gcc.ttp’), the preprocessor (‘gcc-cpp.ttp’), the main body (‘gcc-cc1.ttp’), the assembler (‘gcc-as.ttp’) and the linker (‘gcc-ld.ttp’). It also contains the following support programs:

‘gcc-ar.ttp’

is the object library maintainer.

‘gdb.ttp’ is the GNU debugger modified for the Atari ST. John Dunning did the port to the

Atari. Since then Jwahar Bammi has extensively hacked it.

‘sym-ld.ttp’

creates the symbol file needed with gdb.

‘gcc-nm.ttp’

prints the symbols of a GNU archive or object file.

‘cnm.ttp’ prints the symbol table of a GEMDOS executable.

‘fixstk.ttp’
‘printstk.ttp’

are used to modify and print the current stack size of an executable.

‘toglclr.ttp’

TOS 1.4 users can toggle the clear above BSS to end of TPA flag for the GEMDOS loader.

‘gnu.g’ is a sample file for the GULAM PD shell. ‘COPYING’ explains your rights and responsibilities as a GNU user.

‘readme’
‘readme.st’

are notes from Jwahar R. Bammi and John R. Dunning.

All the executables should go in ‘\gnu\bin\’ on your gnu disk, or where ever the environment variable GCCEXEC will point at. The executables assume that you’re running an approprate shell, one that will pass command line arguements using either the Mark Williams or Atari conventions. Gulam and Master are two such shells, Gemini a german alternate shareware desktop also works. Other CLI’s may also work. The compiler driver ‘gcc.ttp’ should be in the search PATH of your shell. The next step is to define GCCEXEC. ‘gcc.ttp’ uses this variable to locate the preprocessor, compiler, assembler and the linker. GCCEXEC contains a device/dir/partial-pathname. Assuming you also put the executables in the directory ‘c:\gnu\bin’, GCCEXEC would contain ‘c:\gnu\bin\gcc-’. The value is the same as you would specify in the ‘-B’ option to the compiler driver.

Then you should define a variable called TEMP. During compilation the ouput of the various stages is kept here. The variable must not contain a trailing backslash. If you have enough memory, TEMP should point to a ramdisk.

The next thing to do is to install the libraries. The distributed archive contains the following libraries:

‘README’ The obvious. ‘crt0.o’ is the startup module. ‘gcrt0.o’ is the startup module for profiling runs. Automatically selected by the compiler driver

when you specify the ‘-pg’ option.

‘gnu.olb’
‘gnu16.olb’

are the standard libraries, the usual ‘libc’ on other systems.

‘curses.olb’
‘curses16.olb’

are ports of the BSD curses.

‘gem.olb’
‘gem16.olb’

contain the Atari ST Aes/Vdi/FSM-Gdos bindings.

‘iio.olb’
‘iio16.olb’

contain the integer only ‘printf’ and ‘scanf’ functions.

‘pml.olb’
‘pml16.olb’

are the portable math libraries.

‘widget.olb’
‘widget16.olb’

are a small widget set based on ‘curses’

All these libraries go into a subdirectory described by the environment variable GNULIB. Again this variable must not contain a trailing backslash. Staying with the above example, I’ve set the variable to ‘c:\gnu\lib’. The libraries, which have a 16 in their names were compiled with the ‘-mshort’ option. This makes integers the same size as shorts.

The last bit to install are the header files. They are contained in an archive of their own. The preprocessor now knows about the variable GNUINC. Earlier version had to use the ‘-Iprefix’ option, to get to the header files. According to the above examples, the files would be put in the directory ‘c:\gnu\include’. GNUINC has to be set accordingly.

With earlier versions of GNU CC it was only allowed to put one path into the variables GNULIB and GNUINC. GCC 1.37 and later allows you to put several paths into these variables, which are separated by either a , or a ;. All the mentioned paths are searched in order to locate a specific file. However the startup modules ‘crt0.o’ or ‘gcrt0.o’ are only looked for in the first directory specified in GNULIB. If the preprocessor can’t find a include file in one of the directories specified by GNUINC, it will also search the paths listed in GNULIB.

The programs, which come with the GCC distribution also understand filenames, which use the slash (‘/’) as a separator. If Gulam is your favorite CLI you will stick to the backslashes, since you otherwise lose the feature of command line completition.

If you are using Gulam, you can define ‘aliases’ to reach the executables under more common names.

     alias cc c:\gnu\bin\gcc.ttp
     alias ar c:\gnu\bin\gcc-ar.ttp
     alias as c:\gnu\bin\gcc-as.ttp

…

Now you should be able to say ‘cc foo.c -o foo.ttp’ and the obvious things should happen. If you still have trouble, compare your settings with the ones from the sample file ‘gulam.g’. That should give you the rigth idea.

One additional note to Gulam. ‘crt0.o’ is currently set up to understand the MWC/Atari convention of passing long command lines (execpt it doesn’t look into the _io_vector part). Gulam users should set ‘env_style mw’, if you want to give long args lines to ‘gcc.ttp’.

Memory Requirements

GCC loves memory. A lot. It loves to cons structures. Lots of them. All versions of GCC run in 1 Meg of memory, but to get compile any “real” programs atleast 2 Megs are recommended. In versions before 1.37 the ‘gcc-cc1.ttp’ had 1/2 meg stack, and needs it for compiling large files with optimization turned on. Happily, it doesn’t need all that stack for smaller files, or even big files without the ‘-O’ option, so it should be feasible to make a compiler with a smaller stack (with

‘fixstk.ttp’).

GCC version 1.37 uses another scheme for memory allocation. The programs ‘gcc-cpp.ttp’ and ‘gcc-cc1.ttp’ are setup for _stksize == -1L. This means, that an executable will use all available memory, doing mallocs from internal heap (as opposed to the system heap via Malloc), with SP initially set at the top, and heap starting just above the BSS. So if the compiler runs out of memory, you probably need more memory (or get rid of accessories, tsr’s etc and try).

2 Controlling the Compiler Driver

The GNU C compiler uses a command syntax much like the U**x C compiler. The ‘gcc.ttp’ program accepts options and file names as operands. Multiple single-letter options may not be grouped: ‘-dr’ is very different from ‘-d -r’.

When you invoke GNU CC, it normally does preprocessing, compilation, assembly and linking. File names which end in ‘.c’ are taken as C source to be preprocessed and compiled; file names ending in ‘.i’ are taken as preprocessor output to be compiled; compiler output files plus any input files with names ending in ‘.s’ are assembled; then the resulting object files, plus any other input files, are linked together to produce an executable.

Command options allow you to stop this process at an intermediate stage.

For example, the ‘-c’ option says not to run the linker. Then the output consists of object files output by the assembler.

Other command options are passed on to one stage of processing. Some options control the preprocessor and others the compiler itself. Yet other options control the assembler and linker; these are not documented here, but you rarely need to use any of them.

Here are the options to control the overall compilation process, including those that say whether to link, whether to assemble, and so on.

‘-o file’ Place output in file file. This applies regardless to whatever sort of output is being produced, whether it be an executable file, an object file, an assembler file or preprocessed C code. If ‘-o’ is not specified, the default is to put an executable file in ‘a.out’, the object file ‘source.c’ in ‘source.o’, an assembler file in ‘source.s’, and preprocessed C on standard output.

‘-c’ Compile or assemble the source files, but do not link. Produce object files with names made by replacing ‘.c’ or ‘.s’ with ‘.o’ at the end of the input file names. Do nothing at all for object files specified as input.

‘-S’ Compile into assembler code but do not assemble. The assembler output file name is made by replacing ‘.c’ with ‘.s’ at the end of the input file name. Do nothing at all for assembler source files or object files specified as input.

‘-E’ Run only the C preprocessor. Preprocess all the C source files specified and output the results to standard output.

‘-v’ Compiler driver program prints the commands it executes as it runs the preprocessor, compiler proper, assembler and linker. Some of these are directed to print their own version numbers.

‘-s’ The executable is stripped from the DRI compatible symbol table. Certain symbolic debuggers like ‘sid.prg’ work with this symbol table. Also the programs ‘printstk.ttp’ and ‘fixstk.ttp’ (See see Chapter 5 [The Utilities], page 36, for more info) lookup the symbol ‘_stksize’ in this table.

‘-Bprefix’ Compiler driver program tries prefix as a prefix for each program it tries to run. These

programs are ‘gcc-cpp.ttp’, ‘gcc-cc1.ttp’, ‘gcc-as.ttp’ and ‘gcc-ld.ttp’.

For each subprogram to be run, the compiler driver first tries the ‘-B’ prefix, if any. If that name is not found, or if ‘-B’ was not specified, the driver tries two standard prefixes, which are ‘/usr/lib/gcc-’ and ‘/usr/local/lib/gcc-’. If neither of those results in a file name that is found, the unmodified program name is searched for using the directories specified in your ‘PATH’ environment variable. The run-time support file ‘gnu.olb’ is also searched for using the ‘-B’ prefix, if needed. If it is not found there, the two standard prefixes above are tried, and that is all. The file is left out of the link if it is not found by those means. Most of the time, on most machines, you can do without it. You can get a similar result from the environment variable GCCEXEC. If it is defined, its value is used as a prefix in the same way. If both the ‘-B’ option and the GCCEXEC variable are present, the ‘-B’ option is used first and the environment variable value second.

These options control the details of C compilation itself.

‘-ansi’ Support all ANSI standard C programs.

This turns off certain features of GNU C that are incompatible with ANSI C, such as the asm, inline and typeof keywords, and predefined macros such as unix and vax that identify the type of system you are using. It also enables the undesirable and rarely used ANSI trigraph feature. The ‘-ansi’ option does not cause non-ANSI programs to be rejected gratuitously. For that, ‘-pedantic’ is required in addition to ‘-ansi’. The macro __STRICT_ANSI__is predefined when the ‘-ansi’ option is used. Some header files may notice this macro and refrain from declaring certain functions or defining certain macros that the ANSI standard doesn’t call for; this is to avoid interfering with any programs that might use these names for other things.

‘-traditional’

Attempt to support some aspects of traditional C compilers. Specifically:

• All extern declarations take effect globally even if they are written inside of a function definition. This includes implicit declarations of functions.

• The keywords typeof, inline, signed, const and volatile are not recognized.

• Comparisons between pointers and integers are always allowed.

  • Integer types unsigned short and unsigned char promote to unsigned int.

• Out-of-range floating point literals are not an error.

• All automatic variables not declared register are preserved by longjmp. Ordinarily, GNU C follows ANSI C: automatic variables not declared volatile may be clobbered.

• In the preprocessor, comments convert to nothing at all, rather than to a space. This allows traditional token concatenation.

• In the preprocessor, macro arguments are recognized within string constants in a macro definition (and their values are stringified, though without additional quote marks, when they appear in such a context). The preprocessor always considers a string constant to end at a newline.

• The predefined macro __STDC__is not defined when you use ‘-traditional’, but __GNUC__is (since the GNU extensions which __GNUC__indicates are not affected by ‘-traditional’). If you need to write header files that work differently depending on whether ‘-traditional’ is in use, by testing both of these predefined macros you can distinguish four situations: GNU C, traditional GNU C, other ANSI C compilers, and other old C compilers.

‘-O’ Optimize. Optimizing compilation takes somewhat more time, and a lot more memory for a large function. Without ‘-O’, the compiler’s goal is to reduce the cost of compilation and to make debugging produce the expected results. Statements are independent: if you stop the program with a breakpoint between statements, you can then assign a new value to any variable or change the program counter to any other statement in the function and get exactly the results you would expect from the source code. Without ‘-O’, only variables declared register are allocated in registers. The resulting compiled code is a little worse than produced by PCC without ‘-O’. With ‘-O’, the compiler tries to reduce code size and execution time.Some of the ‘-f’ options described below turn specific kinds of optimization on or off.

‘-g’ Produce debugging information in the operating system’s native format (DBX). GDB also can work with this debugging information. Unlike most other C compilers, GNU CC allows you to use ‘-g’ with ‘-O’. The shortcuts taken by optimized code may occasionally produce surprising results: some variables you declared may not exist at all; flow of control may briefly move where you did not expect it; some statements may not be executed because they compute constant results or their values were already at hand; some statements may execute in different places because they were moved out of loops. Nevertheless it proves possible to debug optimized output. This makes it reasonable to use the optimizer for programs that might have bugs.

‘-gg’ This option is no long support. Do not use it.

‘-w’ Inhibit all warning messages.

‘-W’ Print extra warning messages for these events:

• An automatic variable is used without first being initialized.

These warnings are possible only in optimizing compilation, because they require data flow information that is computed only when optimizing. They occur only for variables that are candidates for register allocation. Therefore, they do not occur for a variable that is declared volatile, or whose address is taken, or whose size is other than 1, 2, 4 or 8 bytes. Also, they do not occur for structures, unions or arrays, even when they are in registers.

Note that there may be no warning about a variable that is used only to compute a value that itself is never used, because such computations may be deleted by the flow analysis pass before the warnings are printed.

These warnings are made optional because GNU CC is not smart enough to see all the reasons why the code might be correct despite appearing to have an error. Here is one example of how this can happen:

     {
       int x;
       switch (y)
         {
         case 1: x = 1;
           break;
         case 2: x = 4;
           break;
         case 3: x = 5;
         }
       foo (x);
     }

If the value of y is always 1, 2 or 3, then x is always initialized, but GNU CC doesn’t know this. Here is another common case:

     {
       int save_y;
       if (change_y) save_y = y, y = new_y;

…

       if (change_y) y = save_y;
     }

This has no bug because save_y is used only if it is set.

Some spurious warnings can be avoided if you declare as volatile all the functions you use that never return.

• A nonvolatile automatic variable might be changed by a call to longjmp. These warnings as well are possible only in optimizing compilation. The compiler sees only the calls to setjmp. It cannot know where longjmp will be called; in fact, a signal handler could call it at any point in the code. As a result, you may get a warning even when there is in fact no problem because longjmp cannot in fact be called at the place which would cause a problem.

• A function can return either with or without a value. (Falling off the end of the function body is considered returning without a value.) For example, this function would inspire such a warning:

         foo (a)
         {
           if (a > 0)
             return a;
         }

Spurious warnings can occur because GNU CC does not realize that certain functions (including abort and longjmp) will never return.

• An expression-statement contains no side effects. In the future, other useful warnings may also be enabled by this option.

‘-Wimplicit’

Warn whenever a function is implicitly declared.

‘-Wreturn-type’

Warn whenever a function is defined with a return-type that defaults to int. Also warn about any return statement with no return-value in a function whose return-type is

not void.
‘-Wunused’

Warn whenever a local variable is unused aside from its declaration, and whenever a function is declared static but never defined.

‘-Wcomment’

Warn whenever a comment-start sequence ‘/*’ appears in a comment.

‘-Wall’ All of the above ‘-W’ options combined.

‘-Wwrite-strings’

Give string constants the type const char[length] so that copying the address of one into a non-const char * pointer will get a warning. These warnings will help you find at compile time code that can try to write into a string constant, but only if you have been very careful about using const in declarations and prototypes. Otherwise, it will just be a nuisance; this is why we did not make ‘-Wall’ request these warnings.

‘-p’ Generate extra code to write profile information suitable for the analysis program prof. This is useless on the atariST. Use -pg instead.

‘-pg’ Generate extra code to write profile information suitable for the analysis program

gprof.

‘-llibrary’ Search a standard list of directories for a library named library, which is actually a file named ‘$GNULIB\library.olb’. The linker uses this file as if it had been specified precisely by name. The directories searched include several standard system directories plus any that you specify with ‘-L’. Normally the files found this way are library files—archive files whose members are object files. The linker handles an archive file by scanning through it for members which define symbols that have so far been referenced but not defined. But if the file that is found is an ordinary object file, it is linked in the usual fashion. The only difference between using an ‘-l’ option and specifying a file name is that ‘-l’ searches several directories.

‘-Ldir’ Add directory dir to the list of directories to be searched for ‘-l’.

‘-nostdlib’

Don’t use the standard system libraries and startup files when linking. Only the files you specify (plus ‘gnulib’) will be passed to the linker.

‘-mmachinespec’ Machine-dependent option specifying something about the type of target machine. These options are defined by the macro TARGET_SWITCHES in the machine description. The default for the options is also defined by that macro, which enables you to change the defaults. These are the ‘-m’ options defined in the 68000 machine description:

‘-m68000’
‘-mc68000’

Generate output for a 68000. This is the default on the Atari ST.

‘-m68020’
‘-mc68020’

Generate output for a 68020 (rather than a 68000).

‘-m68881’ Generate output containing 68881 instructions for floating point.

‘-msoft-float’

Generate output containing library calls for floating point. This is the default on the Atari ST.

‘-mshort’ Consider type int to be 16 bits wide, like short int and causes the macro __MSHORT__to be defined. Using this option also causes the library ‘library16.olb’ to be linked. (Also see Section 3.2 [Predefined Macros], page 24, for more info)

‘-mint’ Causes gcc to link with MiNT libraries instead of the standard libraries. It also cause the macro __MINT__to be defined.

‘-mnobitfield’

Do not use the bit-field instructions. ‘-m68000’ implies ‘-mnobitfield’. This is the default on the Atari ST.

‘-mbitfield’

Do use the bit-field instructions. ‘-m68020’ implies ‘-mbitfield’.

‘-mrtd’ Use a different function-calling convention, in which functions that take a fixed number of arguments return with the rtd instruction, which pops their arguments while returning. This saves one instruction in the caller since there is no need to pop the arguments there. This calling convention is incompatible with the one normally used on U**x, so you cannot use it if you need to call libraries compiled with the U**x compiler. Also, you must provide function prototypes for all functions that take variable numbers of arguments (including printf); otherwise incorrect code will be generated for calls to those functions. In addition, seriously incorrect code will result if you call a function with too many arguments. (Normally, extra arguments are harmlessly ignored.) The rtd instruction is supported by the 68010 and 68020 processors, but not by the 68000.

‘-fflag’ Specify machine-independent flags. Most flags have both positive and negative forms; the negative form of ‘-ffoo’ would be ‘-fno-foo’. In the table below, only one of the forms is listed—the one which is not the default. You can figure out the other form by either removing ‘no-’ or adding it.

‘-ffloat-store’

Do not store floating-point variables in registers. This prevents undesirable excess precision on machines such as the 68000 where the floating registers (of the 68881) keep more precision than a double is supposed to have. For most programs, the excess precision does only good, but a few programs rely on the precise definition of IEEE floating point. Use ‘-ffloat-store’ for such programs.

‘-fno-asm’

Do not recognize asm, inline or typeof as a keyword. These words may then be used as identifiers.

‘-fno-defer-pop’

Always pop the arguments to each function call as soon as that function returns. Normally the compiler (when optimizing) lets arguments accumulate on the stack for several function calls and pops them all at once.

‘-fstrength-reduce’

Perform the optimizations of loop strength reduction and elimination of

iteration variables.

‘-fcombine-regs’

Allow the combine pass to combine an instruction that copies one register into another. This might or might not produce better code when used in addition to ‘-O’. I am interested in hearing about the difference this makes.

‘-fforce-mem’

Force memory operands to be copied into registers before doing arithmetic on them. This may produce better code by making all memory references potential common subexpressions. When they are not common subexpressions, instruction combination should eliminate the separate register-load. I am interested in hearing about the difference this makes.

‘-fforce-addr’

Force memory address constants to be copied into registers before doing arithmetic on them. This may produce better code just as ‘-fforce-mem’ may.

‘-fomit-frame-pointer’

Don’t keep the frame pointer in a register for functions that don’t need one. This avoids the instructions to save, set up and restore frame pointers; it also makes an extra register available in many functions. It also makes debugging impossible. On some machines, such as the Vax, this flag has no effect, because the standard calling sequence automatically handles the frame pointer and nothing is saved by pretending it doesn’t exist. The machine-description macro FRAME_POINTER_REQUIRED controls whether a target machine supports this flag.

‘-finline-functions’

Integrate all simple functions into their callers. The compiler heuristically decides which functions are simple enough to be worth integrating in this way. If all calls to a given function are integrated, and the function is declared static, then the function is normally not output as assembler code in its own right.

‘-fkeep-inline-functions’

Even if all calls to a given function are integrated, and the function is declared static, nevertheless output a separate run-time callable version of the function.

‘-fwritable-strings’

Store string constants in the writable data segment and don’t uniquize them. This is for compatibility with old programs which assume they can write into string constants. Writing into string constants is a very bad idea; “constants” should be constant.

‘-fcond-mismatch’

Allow conditional expressions with mismatched types in the second and third arguments. The value of such an expression is void.

‘-fno-function-cse’

Do not put function addresses in registers; make each instruction that calls a constant function contain the function’s address explicitly. This option results in less efficient code, but some strange hacks that alter the assembler output may be confused by the optimizations performed when this option is not used.

‘-fvolatile’

Consider all memory references through pointers to be volatile.

‘-fshared-data’

Requests that the data and non-const variables of this compilation be shared data rather than private data. The distinction makes sense only on certain operating systems, where shared data is shared between processes running the same program, while private data exists in one copy per process.

‘-funsigned-char’

Let the type char be the unsigned, like unsigned char. Each kind of machine has a default for what char should be. It is either like unsigned char by default or like signed char by default. (Actually, at present, the default is always signed.) The type char is always a distinct type from either signed char or unsigned char, even though its behavior is always just like one of those two. Note that this is equivalent to ‘-fno-signed-char’, which is the negative

form of ‘-fsigned-char’.
‘-fsigned-char’

Let the type char be signed, like signed char. Note that this is equivalent to ‘-fno-unsigned-char’, which is the negative

form of ‘-funsigned-char’.
‘-ffixed-reg’

Treat the register named reg as a fixed register; generated code should never refer to it (except perhaps as a stack pointer, frame pointer or in some other fixed role). reg must be the name of a register. The register names accepted are machine-specific and are defined in the REGISTER_NAMES macro in the machine description macro file.

This flag does not have a negative form, because it specifies a three-way choice.

‘-fcall-used-reg’

Treat the register named reg as an allocatable register that is clobbered by function calls. It may be allocated for temporaries or variables that do not live across a call. Functions compiled this way will not save and restore the register reg. Use of this flag for a register that has a fixed pervasive role in the machine’s execution model, such as the stack pointer or frame pointer, will produce disastrous results. This flag does not have a negative form, because it specifies a three-way choice.

‘-fcall-saved-reg’

Treat the register named reg as an allocatable register saved by functions. It may be allocated even for temporaries or variables that live across a call. Functions compiled this way will save and restore the register reg if they use it. Use of this flag for a register that has a fixed pervasive role in the machine’s execution model, such as the stack pointer or frame pointer, will produce disastrous results. A different sort of disaster will result from the use of this flag for a register in which function values may be returned. This flag does not have a negative form, because it specifies a three-way choice.

‘-pedantic’

Issue all the warnings demanded by strict ANSI standard C; reject all programs that use forbidden extensions. Valid ANSI standard C programs should compile properly with or without this option (though a rare few will require ‘-ansi’). However, without this option, certain GNU extensions and traditional C features are supported as well. With this option, they are rejected. There is no reason to use this option; it exists only to satisfy pedants.

These options control the C preprocessor, which is run on each C source file before actual compilation. If you use the ‘-E’ option, nothing is done except C preprocessing. Some of these options make sense only together with ‘-E’ because they request preprocessor output that is not suitable for actual compilation.

‘-C’ Tell the preprocessor not to discard comments. Used with the ‘-E’ option. ‘-Idir’ Search directory dir for include files.

‘-I-’ Any directories specified with ‘-I’ options before the ‘-I-’ option are searched only for the case of ‘#include "file"’; they are not searched for ‘#include <file>’. If additional directories are specified with ‘-I’ options after the ‘-I-’, these directories are searched for all ‘#include’ directives. (Ordinarily all ‘-I’ directories are used this way.) In addition, the ‘-I-’ option inhibits the use of the current directory as the first search directory for ‘#include "file"’. Therefore, the current directory is searched only if it is requested explicitly with ‘-I.’. Specifying both ‘-I-’ and ‘-I.’ allows you to control precisely which directories are searched before the current one and which are searched after.

‘-nostdinc’

Do not search the standard system directories for header files. Only the directories you have specified with ‘-I’ options (and the current directory, if appropriate) are searched. Between ‘-nostdinc’ and ‘-I-’, you can eliminate all directories from the search path except those you specify.

‘-M’ Tell the preprocessor to output a rule suitable for make describing the dependencies of each source file. For each source file, the preprocessor outputs one make-rule whose target is the object file name for that source file and whose dependencies are all the files ‘#include’d in it. This rule may be a single line or may be continued with ‘\’-newline if it is long. ‘-M’ implies ‘-E’.

‘-MM’ Like ‘-M’ but the output mentions only the user-header files included with ‘#include "file"’. System header files included with ‘#include <file>’ are omitted. ‘-MM’ implies ‘-E’.

‘-Dmacro’ Define macro macro with the empty string as its definition. ‘-Dmacro=defn’

Define macro macro as defn.

‘-Umacro’ Undefine macro macro. ‘-T’ Support ANSI C trigraphs. You don’t want to know about this brain-damage. The

‘-ansi’ option also has this effect.

3 The Preprocessor

3.1 Invoking the C Preprocessor

Most often when you use the C preprocessor you will not have to invoke it explicitly: the C compiler will do so automatically. However, the preprocessor is sometimes useful individually.

The C preprocessor expects two file names as arguments, infile and outfile. The preprocessor reads infile together with any other files it specifies with ‘#include’. All the output generated by the combined input files is written in outfile.

Either infile or outfile may be ‘-’, which as infile recogly the user-header fileøl2stem direis wrined i!ou y bs+e user ofile info output.

����

�-ˇ tnctiotput.